# Bridging Speech Emotion Recognition and Personality: Dataset and Temporal Interaction Condition Network

### v1

@@ IV SER with Ground-Truth Personality Traits
Given an utterance $U\in\{U_{1},U_{2},...,U_{n}\}$, we incorporate HuBERT [41], which consists of 7 convolutional layers followed by 12 Transformer layers as acoustic feature extractor. This model is pretrained in a self-supervised learning manner and can capture both acoustic features and content information from the input utterance. The acoustic features learned from HuBERT are represented as $A^{e}\in\mathbb{R}^{t\times d}$, where $t$ is the number of temporal frames and $d$ is the hidden dimension of the Transformer layers. Previous works have validated that emotion perception requires both acoustic features and linguistic content. Therefore, we incorporate an ASR component to learn the linguistic information and thus ensure promising performance for SER [22, 42]. We feed $A^{e}$ into a linear layer and apply the connectionist temporal classification (CTC) loss function [43]:
$\mathcal{L}_{\text{ASR}}=\text{CTC}(y,\hat{y}),$ (1)
@@
$\mathcal{L}=(1-\lambda)\mathcal{L}_{\text{SER}}+\lambda\mathcal{L}_{\text{ASR}},$ (3)
@@
Inspired by the correlation between personality and emotion expression in Table II, we first introduce a linear layer to project the personality features $P$ from each Big Five traits $t\in\mathbb{\{}{OP,CO,EX,AG,NE}\}$. The pooled acoustic features $A^{e}_{\text{pooled}}$ and the personality features $P$ are then concatenated to form a combined feature $E\in\mathbb{R}^{d+e}$ for SER. We also provide the comparison of using all personality features as additional information in the experiment section.
@@ V Personality Recognition and SER with predicted Personality Traits
In real-world dialogue system applications, users do not typically provide explicit information about their personality traits. Therefore, instead of relying on ground-truth personality traits, we explore the use of predicted personality traits for improving SER. We first conduct comprehensive PR experiments using PA-IEMOCAP, evaluating recognition performance at both the utterance and conversation levels. Given the predicted personality, our framework enables dialogue systems to infer personality traits directly from speech input. Furthermore, we introduce a TICN to capture the temporal dynamics of personality impact on emotional expression and incorporate cross-attention for effective feature fusion.
@@ V-A Utterance-level Personality Recognition
$\mathcal{L}=(1-\gamma)\mathcal{L}_{\text{PR}}+\gamma\mathcal{L}_{\text{ASR}},$ (5)
where $\lambda$ is a weight parameter balancing the contribution of these two tasks.
@@
$\mathcal{L}=(1-\alpha-\beta)\mathcal{L}_{\text{PR}}+\alpha\mathcal{L}_{\text{SER}}+\beta\mathcal{L}_{\text{ASR}},$ (6)
@@ V-C Temporal Interaction Condition Network (TICN)
Self-attention is shown to be effective for SER as emotional expression varies at different temporal segments of an utterance. We hypothesize the impact of personality traits also varies temporally. To model this dynamic effect, we propose a TICN to extract personality traits into finegrained personality features. Specifically, we unsqueeze the personality feature $P$ projected from linear layer as $P^{t}\in\mathbb{R}^{t\times d}$ to align their dimensions with $A^{p}$. We integrate the temporal context information of the input utterance by performing element-wise addition between $A^{p}$ and $P$, which can be denoted as:
$P=0.9\cdot P\oplus 0.1\cdot A^{p}$ (8)
@@
$E^{\mathrm{TICN}}=\text{softmax}\left((P^{t}W^{Q})(AW^{K})^{T}/\sqrt{d_{k}}\right)(AW^{V})$ (9)
This proposed approach enables $A^{p}$ to selectively incorporate information from $P^{t}$ that is most relevant to emotional expressions at each temporal segment.
As shown in the right sub-figure of Fig. 1, the outputs of the cross-attention layer $E^{\mathrm{TICN}}$ are then pooled to obtain fixed-length representations, which are subsequently used for SER. The overall loss function of the SER is the same as Eq. (3).
TABLE III: Effect of ground-truth Big Five traits in arousal and valence recognition. We use “*” to denote statistically significant improvement (p $<$ 0.05) from the baseline.
@@ VI-A Implementation and Setup
We implemented the proposed model using PyTorch and the Huggingface Transformers repository [44]. Our experiments utilize HuBERT-base [41] as the acoustic feature extractor, which was pretrained on 60,000 hours of Libri-Light data [45] and comprises a convolutional feature extractor followed by 12 Transformer encoder layers with 768-dimensional hidden representations. During training, we strategically froze the CNN feature extractor and the first 6 Transformer layers while fine-tuning the remaining 6 layers, employing the AdamW optimizer with a learning rate of 5e-5, a mini-batch size of 2 with gradient accumulation steps of 8. The model was trained with dynamic batch padding to accommodate varying input lengths. For multi-task learning experiments, the weight parameters $\lambda$, $\gamma$, and $\alpha$, $\beta$ were set to 0.1.
We split the PA-IEMOCAP dataset into a training set (sessions 2–4), a validation set (14 conversations from session 1), and a test set (another 14 conversations from session 1), ensuring speaker-independent evaluation. We employed concordance correlation coefficient (CCC) as the primary metric for both PR and SER and Word Error Rate (WER) for the ASR, with model checkpoints saved based on the best CCC performance on the validation set.
@@ VI-B Effect of Ground-Truth Personality Traits for SER
In this study, we explored three approaches to integrate personality information with acoustic features for enhancing SER. The first method, Concat, involves conditioning personality features through a linear layer and concatenating them with acoustic features. The second method, CA, employs a cross-attention mechanism to combine personality and acoustic features, facilitating effective interactions between the two modalities. The third method, TICN-CA, incorporates TICN to model the temporal dynamics of personality traits, which are then integrated with acoustic features via cross-attention (CA). We compare the proposed approaches with baseline model that uses only speech input. The results are presented in Table III.
We observe that concatenating acoustic features with personality embeddings significantly improved valence recognition, particularly when incorporating features learned from OP and AG. This aligns with the correlation analysis, where OP and AG exhibited the strongest correlations with valence. CA facilitates more effective interactions between personality and acoustic features, leading to noticeable improvements. The proposed TICN captures the temporal dynamics of personality traits’ effects on emotional expression, thereby significantly improving valence recognition. According to Table III, the best performance was achieved by using TICN to project all five traits collectively, which were then combined with acoustic features via cross-attention. This approach yielded a CCC of 0.791 for valence recognition. These findings suggest that integrating all personality traits provides a more robust improvement in valence recognition compared to utilizing any single trait individually. However, improvements in arousal recognition remained limited across all personality traits, indicating that personality information show more influential impact in valence rather than arousal in emotional expressions.
@@
TABLE IV: Experimental results of personality recognition for each trait. Results are reported using concordance correlation coefficient (CCC). Utt: Utterance-level experiment. Conv: Conversational-level experiment.
@@ VI-C1 Utterance-level Personality Recognition
The results of the PR experiments are presented in Table IV. We observe that AG gets the best recognition performance, and CO is the most difficult to predict. This indicates that individual utterances contain personality information, and achieved moderate CCC across all Big Five traits. Applying post-processing by averaging predictions across all utterances from the same speaker, we observed consistent improvements for all five traits. The CCC values increased by an average of 0.048, with the most significant improvement observed for AG and OP. This result validates that postprocessing mitigates the variability inherent in utterance-level experiments, and personality traits manifest more consistently across multiple utterances.
@@ VI-C2 Conversational-Level Personality Recognition
As shown in the rightmost column of Table IV, this approach significantly outperformed utterance-level analysis, with CCC improvements exceeding 0.2 across all traits. AG achieved the best performance, with a CCC of 0.843, while OP exhibited the smallest improvement, achieving 0.695 on CCC. These results validate that predicting personality traits should take into account information from the entire conversation, as conversation-level features effectively capture contextual variations. The significant improvement between these two settings underscores the importance of broader conversational context and speaking patterns in revealing critical personality cues unavailable in single utterances.
@@ VI-D Conditioning Predicted Personality Traits for SER
TABLE V: Valence recognition performance when conditioning on predicted or ground-truth personality traits using TICN-CA. We evaluate how using different Big Five traits (left) and different PR approaches (middle) affects valence recognition. Results are reported using concordance correlation coefficient (CCC). We use “*” to denote statistically significant improvement (p $<$ 0.05) from the baseline.
@@
The results of Table III and Fig. 3 (a) have shown that incorporating personality information can benefit valence recognition. In this section, we conduct experiments using predicted personality labels and more sophisticated TICN for enhancing SER. The predicted labels were obtained from the three different personality recognition settings described in the previous section. To integrate the personality information, we employed TICN-CA, which showed the best performance in Table III. Table V summarizes the results.
These results indicate that incorporating predicted personality traits improves valence recognition, with the conversational-level PR approach yielding the highest performance across all settings. When individual traits were incorporated, OP, AG, and NE led to the most notable improvements, consistent with earlier findings. The integration of all five traits predicted by conversational-level PR achieved the best results. In contrast, incorporating traits predicted from single utterance showed limited performance. These results further validate the effectiveness of personality-aware SER and highlight the advantages of leveraging conversational context for more reliable personality trait estimation.
@@ VII Conclusions
These findings validate the efficacy of personality-aware SER systems and highlight their potential for real-world applications where explicit personality data is unavailable. By demonstrating that predicted personality traits can substantively enhance emotion recognition, our approach paves the way for more personalized and emotionally intelligent human-computer interaction systems. This work not only addresses the gap in available emotion and personality annotated datasets but also establishes a foundation for future research into multimodal and context-aware affective computing.
@@ VIII Acknowledgment
This work was supported by JST SPRING (JPMJSP2110), JST Moonshot R&\&&D (JPMJPS2011).

### v2

@@ IV SER with Ground-Truth Personality Traits
Given an utterance $U\in\{U_{1},U_{2},...,U_{n}\}$, we incorporate HuBERT [hsu2021hubert], which consists of 7 convolutional layers followed by 12 Transformer layers as acoustic feature extractor. This model is pretrained in a self-supervised learning manner and can capture both acoustic features and content information from the input utterance. The acoustic features learned from HuBERT are represented as $A^{e}\in\mathbb{R}^{t\times d}$, where $t$ is the number of temporal frames and $d$ is the hidden dimension of the Transformer layers. Previous works have validated that emotion perception requires both acoustic features and linguistic content. Therefore, we incorporate an ASR component to learn the linguistic information and thus ensure promising performance for SER [cai2021speech, gao2024enhancing]. We feed $A^{e}$ into a linear layer and apply the connectionist temporal classification (CTC) loss function [graves2006connectionist]:
[…]
@@
$\mathcal{L}_{\text{SER\_baseline}}=(1-\lambda)\mathcal{L}_{\text{SER}}+\lambda\mathcal{L}_{\text{ASR}},$ (3)
@@
Inspired by the correlation between personality and emotion expression in Table II and III, we first introduce a linear layer to project the personality features $P$ from each Big Five traits $t\in\mathbb{\{}{OP,CO,EX,AG,NE}\}$. The pooled acoustic features $A^{e}_{\text{pooled}}$ and the personality features $P$ are then concatenated to form a combined feature $E\in\mathbb{R}^{d+e}$ for SER. We also provide the comparison of using all personality features as additional information in the experiment section.
@@ V Personality Recognition and SER with predicted Personality Traits
In real-world dialogue system applications, users do not typically provide explicit information about their personality traits. Therefore, instead of relying on ground-truth personality traits, we explore the use of predicted personality traits for improving SER. We first conduct comprehensive PR experiments using PA-IEMOCAP, evaluating recognition performance at both the utterance and conversation levels. Given the predicted personality, our framework enables dialogue systems to infer personality traits directly from speech input. Furthermore, we introduce a TICN to capture the temporal dynamics of personality impact on emotional expression and incorporate cross-attention for effective feature fusion (Fig. 1).
@@ V-A Utterance-level Personality Recognition
$\mathcal{L}_{\text{PR\_baseline}}=(1-\gamma)\mathcal{L}_{\text{PR}}+\gamma\mathcal{L}_{\text{ASR}},$ (5)
where $\gamma$ is a weight parameter balancing the contribution of these two tasks.
@@
$\mathcal{L}_{\text{multitask}}=(1-\alpha-\beta)\mathcal{L}_{\text{PR}}+\alpha\mathcal{L}_{\text{SER}}+\beta\mathcal{L}_{\text{ASR}},$ (6)
@@ V-C Temporal Interaction Condition Network (TICN)
Self-attention is shown to be effective for SER as emotional expression varies at different temporal segments of an utterance. Although the personality of a speaker remains consistent, we hypothesize that its influence on emotion expression varies across different temporal segments of an utterance. To enable the model to capture this dynamic alignment, we replicate $P$ across time steps, not to imply temporal variability in personality itself, but to allow personality information to interact with acoustic features at each segment. Through TICN, this design facilitates the extraction of fine-grained personality-conditioned features that reflect the varying salience of personality traits in shaping emotional expression over time. Specifically, we unsqueeze the personality feature $P$ projected from linear layer as $P^{t}\in\mathbb{R}^{t\times d}$ to align their dimensions with $A^{e}$ (acoustic feature for SER). We integrate the temporal emotion information of the input utterance by performing element-wise addition between $A^{e}$ and $P$, which can be denoted as:
$P=0.9\cdot P\oplus 0.1\cdot A^{e}$ (8)
@@
$E^{\mathrm{ticn}}=\text{softmax}\left((P^{t}W^{Q})(A^{e}W^{K})^{T}/\sqrt{d_{k}}\right)(A^{e}W^{V})$ (9)
Here, $P^{t}$ serve as the query, while $A^{e}$ act as both the key and the value, with the number of attention heads set to 4. We also examined the reverse configuration, in which $P^{t}$ acted as key/value and $A^{e}$ as query; however, this design yielded unsatisfactory results and was not further considered. This proposed approach enables $A^{e}$ to selectively incorporate information from $P^{t}$ that is most relevant to emotional expressions at each temporal segment.
As shown in the right sub-figure of Fig. 1, the outputs of the cross-attention layer $E^{\mathrm{ticn}}\in\mathbb{R}^{t\times d}$ are then pooled to obtain fixed-length representations, which are subsequently used for SER. The overall loss function of SER follows the same formulation as in Eq. (3).
TABLE IV: Effect of ground-truth Big Five traits in arousal and valence recognition. We use “*” to denote statistically significant improvement (p << 0.05) from the baseline.
@@ VI-A Implementation and Setup
We implemented the proposed model using PyTorch and the Huggingface Transformers repository [wolf2020Transformers]. Our experiments utilize HuBERT-base [hsu2021hubert] as the acoustic feature extractor, which was pretrained on 60,000 hours of Libri-Light data [kahn2020libri] and comprises a convolutional feature extractor followed by 12 Transformer encoder layers with 768-dimensional hidden representations. For the linear layers in TICN, the output dimension was fixed at 768 to ensure consistency between $P^{t}$ and $A^{e}$. During training for both SER and PR tasks, the CNN feature extractor together with the first six Transformer layers were frozen, while the remaining six layers were fine-tuned. Optimization was performed using AdamW with a learning rate of $5\times 10^{-5}$, a mini-batch size of 2, and gradient accumulation over 8 mini-batches. The model was trained with dynamic batch padding to accommodate varying input lengths. For multi-task learning experiments, the weight parameters $\lambda$, $\gamma$, $\alpha$, and $\beta$ were tuned among 1,0.1,0.01{1,0.1,0.01}, with 0.1 yielding the best performance for all of them and thus adopted in our experiments. Each experiment was repeated three times with the default random seed of Huggingface and we report the average results.
We split the PA-IEMOCAP dataset into a training set (sessions 2–4), a validation set (14 conversations from session 1), and a test set (another 14 conversations from session 1), ensuring speaker-independent evaluation for both SER and PR. We employed concordance correlation coefficient (CCC) [lawrence1989concordance] as the primary metric for both PR and SER, with model checkpoints saved based on the best CCC performance (SER) on the validation set.
@@ VI-B Effect of Ground-Truth Personality Traits for SER
In this study, we explored three approaches to integrate personality information with acoustic features for enhancing SER. The first method, Concat, involves conditioning personality features through a linear layer and concatenating them with acoustic features. The second method, CA, employs a cross-attention mechanism to combine personality and acoustic features, facilitating effective interactions between the two modalities. The third method, TICN-CA, incorporates TICN to model the temporal dynamics of personality traits, which are then integrated with acoustic features via cross-attention (CA). We compare the proposed approaches with baseline model that uses only speech input. The results are presented in Table IV.
We observe that concatenating acoustic features with personality embeddings significantly improved valence recognition, particularly when incorporating features learned from OP and AG. This aligns with the correlation analysis, where OP and AG exhibited the strongest correlations with valence. CA facilitates more effective interactions between personality and acoustic features, leading to noticeable improvements. The proposed TICN captures the temporal dynamics of personality traits’ effects on emotional expression, thereby significantly improving valence recognition. According to Table IV, the best performance was achieved by using TICN to project all five traits collectively, which were then combined with acoustic features via cross-attention. This approach yielded a CCC of 0.791 for valence recognition. The proposed model introduces only a few additional linear layers, one Transformer block, and a cross-attention mechanism, resulting in a relatively small parameter increase of about 12.1% (106.8M vs. 95.4M in the baseline). These findings suggest that integrating all personality traits provides a more robust improvement in valence recognition compared to utilizing any single trait individually. However, improvements in arousal recognition remained limited across all personality traits, indicating that personality information show more influential impact in valence rather than arousal in emotional expressions.
@@
TABLE V: Experimental results of personality recognition for each trait. Results are reported using concordance correlation coefficient (CCC). Utt: Utterance-level experiment. Conv: Conversational-level experiment.
@@ VI-C1 Utterance-level Personality Recognition
The results of the PR experiments are presented in Table V. We observe that AG gets the best recognition performance, and CO is the most difficult to predict. This indicates that individual utterances contain personality information, and achieved moderate CCC across all Big Five traits. Applying post-processing by averaging predictions across all utterances from the same speaker, we observed consistent improvements for all five traits. The CCC values increased by an average of 0.048, with the most significant improvement observed for AG and OP. This result validates that postprocessing mitigates the variability inherent in utterance-level experiments, and personality traits manifest more consistently across multiple utterances.
@@ VI-C2 Conversational-Level Personality Recognition
As shown in the rightmost column of Table V, this approach significantly outperformed utterance-level analysis, with CCC improvements exceeding 0.2 across all traits. AG achieves the best performance, with a CCC of 0.843, while OP exhibites the smallest improvement, achieving 0.695 on CCC.
These results validate that predicting personality traits should take into account information from the entire conversation, as conversation-level features effectively capture contextual variations. The significant improvement between these two settings underscores the importance of broader conversational context and speaking patterns, suggesting that, much like human perception, computational models also require sufficient conversational evidence to form a reliable impression of a speaker’s personality.
@@ VI-D Conditioning Predicted Personality Traits for SER
TABLE VI: Valence recognition performance when conditioning on predicted or ground-truth personality traits using TICN-CA. We evaluate how using different Big Five traits (left) and different PR approaches (middle) affects valence recognition. Results are reported using concordance correlation coefficient (CCC). We use “*” to denote statistically significant improvement (p << 0.05) from the baseline.
@@
The results of Table IV and Fig. 3 (a) have shown that incorporating personality information can benefit valence recognition. In this section, we conduct experiments using predicted personality labels and more sophisticated TICN for enhancing SER. The predicted labels were obtained from the three different personality recognition settings described in the previous section. To integrate the personality information, we employed TICN-CA, which showed the best performance in Table IV. Table VI summarizes the results.
These results indicate that incorporating predicted personality traits improves valence recognition, with the conversational-level PR approach yielding the highest performance across all settings. When individual traits were incorporated, OP, AG, and NE led to the most notable improvements, consistent with earlier findings. The integration of all five traits predicted by conversational-level PR achieved the best results. In contrast, incorporating traits predicted from single utterance showed limited performance. These results further validate the effectiveness of personality-aware SER and highlight the advantages of leveraging conversational context for more reliable personality trait estimation. Notably, the introduction of the conversational PR module increases the parameter count substantially (from 95.4M in the baseline SER model to 333.9M in total), which should be considered as the trade-off for the performance gain.
@@
TABLE VII: Different levels of Gaussian noise injected into the personality labels during inference. Results are reported using concordance correlation coefficient (CCC).
Gaussian noise Valence (CCC)
0.0 (GT) 0.785
0.1 0.782
0.2 0.780
0.4 0.778
0.8 0.743
1.2 0.697
1.6 0.657
In practical applications, a valence recognition model may be trained with ground-truth personality information but must rely on automatically predicted traits during inference, whose accuracy gradually improves as the dialogue progresses. To examine robustness under such conditions, we injected zero-mean Gaussian noise with varying standard deviations into the personality labels at inference and evaluated its impact on valence recognition. As shown in Table VII, the model remains stable when the noise standard deviation is below 0.8, achieving results comparable to those obtained with ground-truth personality labels (0.785 vs. 0.743). This finding indicates that even coarse or approximate personality estimates are sufficient to provide meaningful benefits for valence recognition. Performance declines more noticeably under larger perturbations (e.g., CCC = 0.657 at $\sigma=1.6$), yet the results demonstrate that the model maintains considerable robustness to moderate levels of personality recognition error.
VI-E Comparasion with existing approaches
To facilitate comparison with prior studies, we also report the SER results under the leave-one-session-out (LOSO) data split.
TABLE VIII: Comparison with existing valence recognition approaches.
Approach Input Year CCC
Srinivasan [srinivasan2022representation] Speech 2022 0.582
Wagner [wagner2023dawn] Speech 2023 0.478
Ispas [ispas2023multi] Speech & Text 2023 0.744
Vlasenko [vlasenko2024comparing] Speech 2024 0.683
Zhou [zhou2024learning] Speech 2024 0.674
Messaoudi [messaoudi2024modeling] Text 2024 0.724
Baseline (w/o traits) Speech 2025 0.681
Proposed (Predicted traits) Speech 2025 0.736
Proposed (GT traits) Speech 2025 0.766
The results in Table VIII demonstrate the rapid progress of valence recognition systems over recent years. Notably, text-based models (e.g., [messaoudi2024modeling]) generally outperform speech-based counterparts, suggesting that lexical cues provide highly informative signals for valence estimation. Furthermore, multimodal approaches that jointly leverage speech and text (e.g., [ispas2023multi]) have achieved the highest reported CCC scores in prior studies, highlighting the complementary nature of linguistic and acoustic information in capturing valence.
As shown in Table VIII, our baseline system achieved performance comparable to other speech-based models by incorporating ASR as an auxiliary task. Incorporating ground-truth (GT) personality traits yielded the best performance, with a CCC of 0.766. When conditioning on predicted personality labels, our model also surpassed both existing speech- and text-based approaches and achieved comparable performance with multimodal system using both speech and text. It is important to note that text input was used only during training for ASR; during inference, the model relied solely on speech. Therefore, the most appropriate comparison is with other speech-based systems.
@@ VII Conclusions
These findings validate the efficacy of personality-aware SER systems and highlight their potential for real-world applications where explicit personality data is unavailable. By demonstrating that predicted personality traits can substantively enhance emotion recognition, our approach paves the way for more personalized and emotionally intelligent human-computer interaction systems. This work not only addresses the gap in available emotion and personality annotated datasets but also establishes a foundation for future research into multimodal and context-aware affective computing. While the personality annotations were derived from acted performances, we also aim to extend our study to more naturalistic data in future work.
@@ VIII Acknowledgment
This work was supported by JST SPRING (JPMJSP2110), and JST Moonshot R&\&D (JPMJPS2011).
[Uncaptioned image] Dr. Yuan Gao received the B.E. degree in school of computer science from Hebei University of Technology, Tianjin, China, in 2019, and the M.S. degree from both Tianjin University, Tianjin, China, and the Japan Advanced Institute of Science and Technology, Ishikawa, Japan, in 2022. He received the Ph.D. degree in Intelligence Science and Technology from Kyoto University, Kyoto, Japan, in 2025. He is currently a Research Scientist with SB Intuitions. His research interests include speech signal processing and multimodal emotion recognition.
[Uncaptioned image] Dr. Hao Shi (Member, IEEE) received the B.E. degree in computer science from Southwest Jiaotong University, Chengdu, China, in 2018, and the M.S. degree in computer science from Tianjin University, Tianjin, China, in 2021. He received the Ph.D. degree in informatics from Kyoto University, Kyoto, Japan, in 2024. From October 2024 to March 2025, he was a researcher at Kyoto University. He is currently a research scientist with SB Intuitions. His research interests include automatic speech recognition and speech enhancement.
[Uncaptioned image] Dr. Yahui Fu received her M.S. degrees from both Tianjin University, Tianjin, China, and the Japan Advanced Institute of Science and Technology, Ishikawa, Japan, in 2021. She received her Ph.D. degree in Intelligence Science and Technology from Kyoto University, Kyoto, Japan, in 2024. She is currently a postdoctoral researcher at Kyoto University. Her research interests include dialogue systems and multimodal emotion recognition.
[Uncaptioned image] Chenhui Chu received his B.S. in software engineering from Chongqing University in 2008, and his M.S. and Ph.D. in Informatics from Kyoto University in 2012 and 2015, respectively. He is currently a associate professor at Kyoto University. His research interests include natural language processing, particularly machine translation and multimodal machine learning.
[Uncaptioned image] Tatsuya Kawahara (Fellow, IEEE) received B.E. in 1987, M.E. in 1989, and Ph.D. in 1995, all in information science, from Kyoto University, Kyoto, Japan. From 1995 to 1996, he was a Visiting Researcher at Bell Laboratories, Murray Hill, NJ, USA. Currently, he is a Professor of School of Informatics, Kyoto University. From 2020 to 2023, he was the Dean of the School. Before that, he was also an Invited Researcher at ATR and NICT. He has published more than 450 academic papers on automatic speech recognition, spoken language processing, and spoken dialogue systems. He has been conducting several projects including open-source speech recognition software Julius, the automatic transcription system deployed in the Japanese Parliament (Diet), and the autonomous android ERICA. Dr. Kawahara received the Commendation for Science and Technology by the Minister of Education, Culture, Sports, Science and Technology (MEXT) in 2012. From 2003 to 2006, he was a member of IEEE SPS Speech Technical Committee. He was a General Chair of IEEE ASRU 2007 and is a General Chair of SIGdial 2024. He also served as a Tutorial Chair of INTERSPEECH 2010, a Local Arrangement Chair of ICASSP 2012, and a General Chair of APSIPA ASC 2020. He was an editorial board member of Elsevier Journal of Computer Speech and Language and IEEE/ACM Transactions on Audio, Speech, and Language Processing. From 2018 to 2021, he was the Editor-in-Chief of APSIPA Transactions on Signal and Information Processing. Dr. Kawahara is the President of APSIPA, the Secretary General of ISCA, and a Fellow of IEEE.